# Learning Manipulation Skills through Robot Chain-of-Thought with Sparse Failure Guidance

### v2

@@ 5.1 Overall Comparison
Our findings indicate distinct performance trends across various tasks. RoboCLIP displays promising results in tasks involving drawer and window manipulation, as well as soccer, but it is most effective in the simplest tasks such as drawer closing. CLIP performs well in tasks related to faucet, drawer and window manipulation, and soccer, indicating that CLIP focuses more on task completion rather than the manipulation process. LIV shows competence primarily in tasks involving drawer closure, window opening, and soccer. Our algorithm RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as closing and opening drawers, windows, and doors. These results underscore the effectiveness of our approach in handling diverse and challenging scenarios.
5.2 Ablation Study
@@
In this section, we conduct ablation studies by removing each component of our algorithm (CoT interpretation, human failure guidance, and self imitation). The results indicate that the most contributing factor to our algorithm’s performance is VLM-based self imitation procedure, followed by human failure guidance. Each component contributes to the overall performance of our algorithm. Additionally, we conduct ablation studies comparing GPT-4V success labels with ground truth success labels, demonstrating the robustness of our GPT-4V labeling. Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s superiority is not solely due to self imitation but also the reward design mechanism.
@@
Question 1: how significant is each component of our algorithm in contributing to the results?
Figure 4: Final policy performance for ablations and our algorithm. The results show that VLM-based self imitation procedure is the most significant contributing factor in our algorithm, followed by human failure guidance.
@@
Our algorithm incorporates several critical components: robot chain-of-thought, sparse human failure guidance, and VLM-based self imitation. To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance. A decline in performance upon the removal of any component demonstrates its essential role in our algorithm.
Figure 4 illustrates the success rates of the final policy performance of our algorithm compared to its ablated versions. The labels for these ablations are: w/oReg (without success experience regularization, i.e., self imitation), w/oHG (without sparse human failure guidance), and w/oCoT (without robot chain-of-thought task decomposition).
The results confirm the critical importance of each component in our algorithm, with VLM-based self imitation being particularly pivotal. Human failure guidance is the second most important component, motivating the agent to consider potential failure cases. Additionally, robot chain-of-thought processing helps the agent better understand the process of manipulation tasks, further contributing to the performance of our algorithm.
Question 2: what is the empirical difference between the labels assigned by GPT-4V and the ground truth labels in the context of success experience collection?
@@
Figure 5: Ablation study for our algorithm with different success labeling methods (GPT-4V labeling or ground truth (GT) labeling). The results demonstrate that the success labels generated by GPT-4V are robust.
@@
Figure 6: Ablation study for all baselines with success experience regularization. The results show that our algorithm performs well not only because of the VLM-based self imitation component but also due to the reward design mechanism.
@@
Our algorithm RoboCoT incorporates GPT-4V to gather successful trajectories for VLM-based self imitation procedure. To validate the robustness of our GPT-4V labeling in collecting success experiences, we conduct experiments comparing GPT-4V labels and ground truth labels. If the performance using GPT-4V labels is similar to that using ground truth labels, it would confirm that GPT-4V is robust in collecting success experiences.
@@
Figure 6 displays the training success rate curves comparing GPT-4V and ground truth (GT) labels for success experience collection. Our findings show that the GPT-4V labels achieves comparable performance to the ground truth across various tasks, demonstrating the robustness of our GPT-4V labels.
Question 3: what is the empirical difference between our algorithm and all baselines with VLM-based self imitation?
Question 1 shows that VLM-based self imitation component plays the most important role in our algorithm. To evaluate the significance of the reward design mechanism, we conduct experiments applying VLM-based self imitation procedure to all baselines. If the results show that our algorithm still outperforms other baselines with self imitation procedure, then it proves that our algorithm’s superiority is due not only to the VLM-based self imitation procedure but also to the reward design mechanism.
Figure 6 shows the ablation results comparing our algorithm with other state-of-the-art baselines that include VLM-based self imitation procedure. The results demonstrate that our algorithm outperforms other baselines on almost all tasks. This indicates that our algorithm’s superiority is due not only to the VLM-based self imitation component but also to the reward design mechanism.
6 Conclusion
In this study, we present a novel algorithm RoboCoT, which defines rewards for skill acquisition. Our algorithm RoboCoT simplifies complex tasks through a robot’s chain-of-thought process with sparse failure guidance. It yields a more fine-grained reward guidance. Through VLM-based self imitation, our algorithm strengthens its policies, thereby improving performance. Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage over established baseline models. This research sets a new standard for mastering manipulation skills in robots, and we hope it can be applied to more robotic tasks in the future.
7 Limitation
Our proposed method, RoboCoT, effectively masters a wide range of manipulation tasks through a robot chain-of-thought with sparse failure guidance. It provides a fine-grained reward guidance for the robot agent, facilitating the training of various automation tools. However, these tasks primarily focus on aligning robot movements with detailed language descriptions. RoboCoT does not yield promising results in tasks that emphasize observational preferences, such as soccer.

### v3

@@ V-A Overall Comparison
Our findings indicate distinct performance trends across various tasks. RoboCLIP displays promising results in the simplest tasks. CLIP performs well in tasks related to faucet, drawer and window manipulation. Because CLIP focuses more on task completion rather than comprehending the robotic manipulation process. LIV shows competence primarily in tasks involving obvious task goal definition. Our RoboCoT consistently outperforms the baselines across all tasks, achieving success rates exceeding $75\%$ in tasks such as drawers, windows, and doors. These results underscore the effectiveness of our RoboCoT in handling diverse and challenging scenarios.
V-B Ablation Study
@@
In this section, we conduct ablation studies by removing each component of our algorithm including without self-imitation procedure, without sparse failure guidance and without robot Chain-of-Thought. The results indicate that each component of our RoboCoT contributes to the overall performance. In addition, we conduct ablation studies comparing GPT-4V with ground truth labels in collecting success experience, demonstrating the robustness of GPT-4V labeling. Finally, experiments with all baselines incorporating VLM-based self imitation procedure reveal that our algorithm’s reward design mechanism is superior compared to other baselines.
@@
Question 1: how significant is each component of our algorithm in contributing to the overall performance?
@@
Our algorithm incorporates several critical components: VLM-based self imitation procedure, sparse human failure guidance, and robot Chain-of-Thought process to define the rewards. To evaluate the significance of each component, we conducted a series of experiments, systematically removing each component to observe its impact on performance. A decline in performance upon the removal of any component demonstrates its essential role in our RoboCoT.
@@
Figure 5: Ablation study for our algorithm with different success experience collection methods (GPT-4V or ground truth (GT) labels). The results demonstrate that the GPT-4V are robust in collecting success experience in a subset of robotic tasks.
@@
Figure 6: Ablation study for all baselines with VLM-based self imitation procedure (w/reg). The results show that the reward design mechanism in our RoboCoT is superior than other baselines.
Figure 4 illustrates the success rates of the final policy performance of our RoboCoT compared to its ablated versions. The labels for these ablations are without self-imitation procedure, without failure guidance and without Chain-of-Thought process.
The results confirm the critical importance of each component in our algorithm. Each component in our RoboCoT contributes to the overall performance. In particular, VLM-based self imitation plays an important role in our RoboCoT. For better evaluating the importance of reward design mechanism in RoboCoT, we also conduct an ablation study with all baselines incorporating VLM-based self imitation procedure in Question 3.
Question 2: Is GPT-4V robust in the context of success experience collection?
@@
Our algorithm RoboCoT incorporates GPT-4V to collect success experience for VLM-based self imitation procedure. To validate the robustness of our GPT-4V labels in collecting success experience, we conduct experiments comparing GPT-4V labels to ground truth labels.
@@
Figure 5 presents the training success rate curves comparing GPT-4V and ground truth (GT) labels for collecting success experience. Our findings indicate that GPT-4V’s labeling performance declines in part of the tasks. However, in most tasks, the performance is comparable to that of the ground truth labels, demonstrating that GPT-4V labels are robust for a subset of robotic tasks.
Question 3: Is the reward design mechanism superior in our RoboCoT compared to other baselines?
Question 1 shows that VLM-based self imitation component plays an important role in our RoboCoT. To evaluate the significance of the reward design mechanism, we conduct experiments incorporating VLM-based self imitation procedure to all baselines. If the results show that our algorithm still outperforms other baselines with VLM based self imitation procedure, then it proves that our algorithm’s reward design mechanism is superior than other baselines.
Figure 6 shows the empirical results comparing our RoboCoT with other state-of-the-art baselines which also incorporating VLM-based self imitation procedure. The results demonstrate that our algorithm outperforms other baselines on almost all tasks. It indicates that the reward design mechanism in our RoboCoT is superior in skill learning compared to other existing baselines.
VI CONCLUSIONS
In this study, we present a novel skill learning algorithm, i.e., RoboCoT, which provides more fine-grained reward guidance for learning manipulation skills. Our RoboCoT simplifies complex tasks through a robot chain-of-thought process. It motivates the robot to better comprehend the complex tasks for skill learning. Through VLM-based self imitation procedure, our RoboCoT speeds up the learning process, thereby improving its performance. Empirical results confirm the effectiveness of our approach, demonstrating a distinct advantage compared to other state-of-the-art baselines. This research sets a new standard for mastering manipulation skills in robotics. We hope our RoboCoT can be applied into more complex robotic tasks in the future.